# Hallo3: Highly Dynamic and Realistic Portrait Image Animation with Diffusion Transformer Networks

### v2

@@ Abstract
Existing methodologies for animating portrait images face significant challenges, particularly in handling non-frontal perspectives, rendering dynamic objects around the portrait, and generating immersive, realistic backgrounds. In this paper, we introduce the first application of a pretrained transformer-based video generative model that demonstrates strong generalization capabilities and generates highly dynamic, realistic videos for portrait animation, effectively addressing these challenges. The adoption of a new video backbone model makes previous U-Net-based methods for identity maintenance, audio conditioning, and video extrapolation inapplicable. To address this limitation, we design an identity reference network consisting of a causal 3D VAE combined with a stacked series of transformer layers, ensuring consistent facial identity across video sequences. Additionally, we investigate various speech audio conditioning and motion frame mechanisms to enable the generation of continuous video driven by speech audio. Our method is validated through experiments on benchmark and newly proposed wild datasets, demonstrating substantial improvements over prior methods in generating realistic portraits characterized by diverse orientations within dynamic and immersive scenes. Further visualizations and the source code are available at: https://fudan-generative-vision.github.io/hallo3.
@@ 1 Introduction
In recent years, the field of portrait image animation has witnessed rapid advancements. Early methodologies predominantly employed facial landmarks—key points [31, 43, 45] on the face utilized for the localization and representation of critical regions such as the mouth, eyes, eyebrows, nose, and jawline. Additionally, these methods [10, 26, 44, 51] incorporated 3D parametric models, notably the 3D Morphable Model (3DMM) [3], which captures variability in human faces through a statistical shape model integrated with a texture model. However, the application of explicit approaches grounded in intermediate facial representations is constrained by the accuracy of expression and head pose reconstruction, as well as the richness and precision of the resultant expressions. Simultaneously, significant advancements in Generative Adversarial Networks (GANs) and diffusion models have notably benefited portrait image animation. These advancements [7, 17, 35, 38, 40, 48, 51] enhance the high-resolution and high-quality generation of realistic facial details, facilitate generalized character animation, and enable long-term identity preservation. Recent contributions to the field—including Live Portrait [11], which leverages GAN technology for portrait animation with stitching and retargeting control, as well as various end-to-end methods such as VASA-1 [40], EMO [35], and Hallo [39, 8] employing diffusion models—exemplify these advancements.
@@
Recent advancements in diffusion transformer (DiT)-based video generation models [41, 22, 2, 18] have addressed several challenges associated with traditional video generation techniques, including issues of realism, dynamic movement, and subject generalization. In this paper, we present the first application of a pretrained DiT-based video generative model to the task of portrait image animation. The introduction of this new video backbone model renders previous U-Net-based methods for identity maintenance, audio conditioning, and video extrapolation impractical. We tackle these issues from three distinct perspectives. (1) Identity preservation: We employ a 3D VAE in conjunction with a stack of transformer layers as an identity reference network, enabling the embedding and injection of identity information into the denoising latent codes for self-attention. This facilitates accurate representation and long-term preservation of the facial subject’s identity. (2) Speech audio conditioning: We achieve high alignment between speech audio—serving as motion control information—and facial expression dynamics during training, which allows for precise control during inference. We investigate the use of adaptive layer normalization and cross-attention strategies, effectively integrating audio embeddings through the latter. (3) Video extrapolation: Addressing the limitations of the DiT-based model in generating continuous videos, which is constrained to a maximum of several tens of frames, we propose a strategy for long-duration video extrapolation. This approach uses motion frames as conditional information, wherein the final frames of each generated video serve as inputs for subsequent clip generation.
@@ 2 Related Work
Portrait Image Animation. Recent advancements in the domain of portrait image animation have been significantly propelled by innovations in audio-driven techniques. Notable frameworks, such as LipSyncExpert [23] and SadTalker [46], have tackled challenges related to facial synchronization and expression modulation, achieving dynamic lip movements and coherent head motions. Concurrently, DiffTalk [30] and VividTalk [33] have integrated latent diffusion models, enhancing output quality while generalizing across diverse identities without the necessity for extensive fine-tuning. Furthermore, studies such as DreamTalk [19] and EMO [35] underscore the importance of emotional expressiveness by showcasing the integration of audio cues with facial dynamics. AniPortrait [38] and VASA-1 [40] propose methodologies that facilitate the generation of high-fidelity animations, emphasizing temporal consistency along with effective exploitation of static images and audio clips. In addition, recent innovations like LivePortrait [11] and Loopy [14] focus on enhancing computational efficiency while ensuring realism and fluid motion. Furthermore, the works of Hallo [39] and Hallo2 [8] have made significant progress in extending capabilities to facilitate long-duration video synthesis and integrating adjustable semantic inputs, thereby marking a step towards richer and more controllable content generation. Nevertheless, existing facial animation techniques still encounter limitations in addressing extreme facial poses, accommodating background motion in dynamic environments, and incorporating camera movements dictated by textual prompts.
@@
Diffusion-Based Video Generation. Unet-based diffusion model has made notable strides, exemplified by frameworks such as Make-A-Video and MagicVideo [49]. Specifically, Make-A-Video [32] capitalizes on pre-existing Text-to-Image (T2I) models to enhance training efficiency without necessitating paired text-video data, thereby achieving state-of-the-art results across a variety of qualitative and quantitative metrics. Simultaneously, MagicVideo [49] employs an innovative 3D U-Net architecture to operate within a low-dimensional latent space, achieving efficient video synthesis while significantly reducing computational requirements. Building upon these foundational principles, AnimateDiff [12] introduces a motion module that integrates seamlessly with personalized T2I models, allowing for the generation of temporally coherent animations without the need for model-specific adjustments. Additionally, VideoComposer [37] enhances the controllability of video synthesis by incorporating spatial, temporal, and textual conditions, which facilitates improved inter-frame consistency. The development of diffusion models continues with the advent of DiT-based approaches such as CogVideoX [41] and Movie Gen [22]. CogVideoX employs a 3D Variational Autoencoder to improve video fidelity and narrative coherence, whereas Movie Gen establishes a robust foundation for high-quality video generation complemented by advanced editing capabilities. In the present study, we adopt the DiT diffusion formulation to optimize the generalization capabilities of the generated video.
@@ 3.1 Baseline Transformer Diffusion Network
Baseline Network. The CogVideoX model [41] serves as the foundational architecture for our transformer diffusion network, employing a 3D VAE for the compression of video data. In this framework, latent variables are concatenated and reshaped into a sequential format, denoted as $\mathbf{z}_{t}$. Concurrently, the model utilizes the T5 architecture [24] to encode textual inputs into embeddings, represented as $\mathbf{c}_{\text{text}}$. The combined sequences of video latent representations $\mathbf{z}_{t}$ and textual embeddings $\mathbf{c}_{\text{text}}$ are subsequently processed through an expert transformer network. To address discrepancies in feature space between text and video, we implement expert adaptive layer normalization techniques, which facilitate the effective utilization of temporal information and ensure robust alignment between visual and semantic data. Following this integration, a repair mechanism is applied to restore the original latent variable, after which the output is decoded through the 3D causal VAE decoder to reconstruct the video. Furthermore, the incorporation of 3D Rotational Positional Encoding (3D RoPE) [41] enhances the model’s capacity to capture inter-frame relationships across the temporal dimension, thereby establishing long-range dependencies within the video framework.
@@
Within diffusion transformers, four primary conditioning mechanisms are identified: in-context conditioning, cross-attention, adaptive layer normalization (adaLN), and adaLN-zero [20]. Our investigation primarily focuses on cross-attention and adaptive layer normalization (adaLN). Cross-attention enhances the model’s focus on conditional information by treating condition embeddings as keys and values, while latent representations serve as queries. Although adaLN is effective in simpler conditioning scenarios, it may not be optimal for more complex conditional embeddings that incorporate richer semantic details, such as sequential speech audio. Relevant comparative analyses will be elaborated upon in the experimental section.
@@ 3.2 Audio-Driven Transformer Diffusion
Speech Audio Embedding. To extract salient audio features for our proposed model, we utilize the wav2vec framework developed by Schneider et al. [28]. The audio representation is defined as $\mathbf{c}_{\text{audio}}$. Specifically, we concatenate the audio embeddings generated by the final twelve layers of the wav2vec network, resulting in a comprehensive semantic representation capable of capturing various audio hierarchies. This concatenation emphasizes the significance of phonetic elements, such as pronunciation and prosody, which are crucial as driving signals for character generation. To transform the audio embeddings obtained from the pretrained model into frame-specific representations, we apply three successive linear transformation layers, mathematically expressed as: $\mathbf{c}_{\text{audio}}^{(f)}=\mathcal{L}_{3}\left(\mathcal{L}_{2}\left(\mathcal{L}_{1}\left(\mathbf{c}_{\text{audio}}\right)\right)\right)$, where $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ represent the respective linear transformation functions. This systematic approach ensures that the resulting frame-specific representations effectively encapsulate the nuanced audio features essential for the performance of our model.
@@ 3.3 Identity Consistent Transformer Diffusion
To address this issue, we introduce a control condition within the existing diffusion transformer architecture to ensure long-term consistency of facial identity appearance. We explore four strategies (as shown in Figure 4) for appearance conditioning: 1) Face attention, where identity features are encoded by the face encoder and combined with a cross-attention module; 2) Face adaptive norm, which integrates features from the face encoder with an adaptive layer normalization technique; 3) Identity reference network, where identity features are captured by a 3D VAE and combined with some transformer layers; and 4) Face attention and Identity reference network, which encodes identity features using both the face encoder and 3D VAE, combining them with self-attention and cross-attention. Our experiments show that the combination with Face attention and Identity reference net achieves the best performance in our model. For further details, please refer to Section 4.3.
@@ 3.4 Training and Inference
(1) Identity Consistency Phase. In this initial phase, we train the model to generate videos with consistent identity. The parameters of the 3D Variational Autoencoder (VAE) and face image encoder remain fixed, while the parameters of the 3D full attention blocks in both the reference and denoising networks, along with the face attention blocks in the denoising network, are updated during training. The model’s input includes a randomly sampled reference image from the training video, a textual prompt, and the face embedding. The textual prompt is generated using MiniCPM[42], which describes human appearance, actions, and detailed environmental background. The face embedding is extracted via InsightFace[9]. With these inputs, the model generates a video comprising 49 frames.
@@
Figure 5: Illustration of the dataset, including the flow of data processing, data distribution across different metric, and the visualization of some representative portrait images for inference.
@@ 3.5 Dataset
Data Sources The training data used in this work is prepared from three distinct sources to ensure diversity and generalization. Specifically, the sources are: (1) HDTF dataset [47], which contains 8 hours of raw video footage; (2) YouTube data, which consists of 1,200 hours of public raw videos; (3) a large scale movie dataset, which contains film videos of 2,346 hours. Our dataset contains a large scale of human identities and, however, we find that YouTube and movie dataset contains a large amount of noised data. Therefore, we design a data curation pipeline as follows to construct a high-quality and diverse talking dataset, as shown in Figure 5(a).
Video Filtering. During the data pre-processing phase, we implement a series of meticulous filtering steps to ensure the quality and applicability of the dataset. The workflow includes three stages: extraction of single-speaker, motion filter and post-processing. Firstly, we select video of single-speaker. This stage aims to clean the video content to solve camera shot, background noise, etc, using existing tools [21, 4]. After that, we apply several filtering techniques to ensure the quality of head motion, head pose, camera motion, etc [16, 15, 6]. In this stage, we compute all metric scores for each clip, therefore, we can flexibly adjust data screening strategies to satisfy different data requirement of our multiple training stages or strategies. Finally, based on the facial positions detected in previous steps, we crop the videos to a 3:2 aspect ratio to meet the model’s input requirements. We then select a random frame from each video and use InsightFace [25] to encode the face into embeddings, providing essential facial feature information for the model. Additionally, we extract the audio from the videos and encode it into embeddings using Wav2Vec2 model [1], facilitating the incorporation of audio conditions during model training.
Data Statistics. Following the data cleaning and filtering processes, we conducted a detailed analysis of the final dataset to assess its quality and suitability for the intended modeling tasks. Finally, our training data contains about 134 hours of videos, including 6 hours of high-quality data from HDTF dataset, 72 hours of YouTube videos, and 56 hours of movie videos. Figure 5(b) also shows other statistics, such as Lip Sync score (Sync-C and Sync-D), face rotation, face ratio (the ratio of face height to video height).
@@
SadTalker [45] 22.340 203.860 7.885 7.545
DreamTalk [19] 78.147 790.660 6.376 8.364
AniPortrait [38] 26.561 234.666 4.015 10.548
Hallo [39] 20.545 173.497 7.750 7.659
@@
SadTalker [45] 50.015 471.163 6.922 7.921 95.194
DreamTalk [19] 109.011 988.539 5.709 8.743 153.450
AniPortrait [38] 46.915 477.179 2.853 11.709 88.986
Hallo [39] 44.578 377.117 7.191 7.984 78.495
@@
SadTalker [45] 3.845 10.378 2.953 0.220 470.377 313.758
DreamTalk [19] 4.498 11.005 6.958 1.806 835.480 744.177
AniPortrait [38] 1.685 12.025 3.351 1.769 473.173 302.716
Hallo [39] 4.654 10.202 5.268 1.272 394.627 291.052

### v4

# Highly Dynamic and Realistic Portrait Image Animation with Diffusion Transformer Networks

@@ Abstract
Existing methodologies for animating portrait images face significant challenges, particularly in handling non-frontal perspectives, rendering dynamic objects around the portrait, and generating immersive, realistic backgrounds. In this paper, we introduce the first application of a pretrained transformer-based video generative model that demonstrates strong generalization capabilities and generates highly dynamic, realistic videos for portrait animation, effectively addressing these challenges. The adoption of a new video backbone model makes previous U-Net-based methods for identity maintenance, audio conditioning, and video extrapolation inapplicable. To address this limitation, we design an identity reference network consisting of a causal 3D VAE combined with a stacked series of transformer layers, ensuring consistent facial identity across video sequences. Additionally, we investigate various speech audio conditioning and motion frame mechanisms to enable the generation of continuous video driven by speech audio. Our method is validated through experiments on benchmark and newly proposed wild datasets, demonstrating substantial improvements over prior methods in generating realistic portraits characterized by diverse orientations within dynamic and immersive scenes.
@@ 1 Introduction
In recent years, the field of portrait image animation has witnessed rapid advancements. Early methodologies predominantly employed facial landmarks—key points [siarohin2019first, zakharov2020fast, zhang2022sadtalker] on the face utilized for the localization and representation of critical regions such as the mouth, eyes, eyebrows, nose, and jawline. Additionally, these methods [gao2023high, ren2021pirenderer, zhang2023metaportrait, champ2024] incorporated 3D parametric models, notably the 3D Morphable Model (3DMM) [blanz2003face], which captures variability in human faces through a statistical shape model integrated with a texture model. However, the application of explicit approaches grounded in intermediate facial representations is constrained by the accuracy of expression and head pose reconstruction, as well as the richness and precision of the resultant expressions. Simultaneously, significant advancements in Generative Adversarial Networks (GANs) and diffusion models have notably benefited portrait image animation. These advancements [corona2024vlogger, liu2024anitalker, tian2024emo, wei2024aniportrait, xu2024vasa, zhang2024tora, champ2024] enhance the high-resolution and high-quality generation of realistic facial details, facilitate generalized character animation, and enable long-term identity preservation. Recent contributions to the field—including Live Portrait [guo2024liveportrait], which leverages GAN technology for portrait animation with stitching and retargeting control, as well as various end-to-end methods such as VASA-1 [xu2024vasa], EMO [tian2024emo], and Hallo [xu2024hallo, cui2024hallo2] employing diffusion models—exemplify these advancements.
@@
Recent advancements in diffusion transformer-based video generation models [yang2024cogvideox, polyak2024movie, bao2024vidu, liu2024sora] have addressed several challenges associated with traditional video generation techniques, including issues of realism, dynamic movement, and subject generalization. In this paper, we present the first application of a pretrained DiT-based video generative model to the task of portrait image animation. The introduction of this new video backbone model renders previous U-Net-based methods for identity maintenance, audio conditioning, and video extrapolation impractical. We tackle these issues from three distinct perspectives. (1) Identity preservation: We employ a 3D VAE in conjunction with a stack of transformer layers as an identity reference network, enabling the embedding and injection of identity information into the denoising latent codes for self-attention. This facilitates accurate representation and long-term preservation of the facial subject’s identity. (2) Speech audio conditioning: We achieve high alignment between speech audio—serving as motion control information—and facial expression dynamics during training, which allows for precise control during inference. We investigate the use of adaptive layer normalization and cross-attention strategies, effectively integrating audio embeddings through the latter. (3) Video extrapolation: Addressing the limitations of the DiT model in generating contineous videos, which is constrained to a maximum of several tens of frames, we propose a strategy for long-duration video extrapolation. This approach uses motion frames as conditional information, wherein the final frames of each generated video serve as inputs for subsequent clip generation.
@@ 2 Related Work
Portrait Image Animation. Recent advancements in the domain of portrait image animation have been significantly propelled by innovations in audio-driven techniques. Notable frameworks, such as LipSyncExpert [prajwal2020lip] and SadTalker [zhang2023sadtalker], have tackled challenges related to facial synchronization and expression modulation, achieving dynamic lip movements and coherent head motions. Concurrently, DiffTalk [shen2023difftalk] and VividTalk [sun2023vividtalk] have integrated latent diffusion models, enhancing output quality while generalizing across diverse identities without the necessity for extensive fine-tuning. Furthermore, studies such as DreamTalk [ma2023dreamtalk] and EMO [tian2024emo] underscore the importance of emotional expressiveness by showcasing the integration of audio cues with facial dynamics. AniPortrait [wei2024aniportrait] and VASA [xu2024vasa] propose methodologies that facilitate the generation of high-fidelity animations, emphasizing temporal consistency along with effective exploitation of static images and audio clips. In addition, recent innovations like LivePortrait [guo2024liveportrait] and Loopy [jiang2024loopy] focus on enhancing computational efficiency while ensuring realism and fluid motion. Furthermore, the works of Hallo [xu2024hallo] and Hallo2 [cui2024hallo2] have made significant progress in extending capabilities to facilitate long-duration video synthesis and integrating adjustable semantic inputs, thereby marking a step towards richer and more controllable content generation. Nevertheless, existing facial animation techniques still encounter limitations in addressing extreme facial poses, accommodating background motion in dynamic environments, and incorporating camera movements dictated by textual prompts.
@@
Diffusion-Based Video Generation. Unet-based diffusion video generation has made notable strides, exemplified by frameworks such as Make-A-Video and MagicVideo [zhou2022magicvideo]. Specifically, Make-A-Video [singer2022make] capitalizes on pre-existing T2I models to enhance training efficiency without necessitating paired text-video data, thereby achieving state-of-the-art results across a variety of qualitative and quantitative metrics. Simultaneously, MagicVideo [zhou2022magicvideo] employs an innovative 3D U-Net architecture to operate within a low-dimensional latent space, achieving efficient video synthesis while significantly reducing computational requirements. Building upon these foundational principles, AnimateDiff [guo2023animatediff] introduces a motion module that integrates seamlessly with personalized T2I models, allowing for the generation of temporally coherent animations without the need for model-specific adjustments. Additionally, VideoComposer [wang2023videocomposer] enhances the controllability of video synthesis by incorporating spatial, temporal, and textual conditions, which facilitates improved inter-frame consistency. The development of diffusion models continues with the advent of DiT-based approaches such as CogVideoX [yang2024cogvideox] and Movie Gen [polyak2024movie]. CogVideoX employs a 3D Variational Autoencoder to improve video fidelity and narrative coherence, whereas Movie Gen establishes a robust foundation for high-quality video generation complemented by advanced editing capabilities. In the present study, we adopt the DiT diffusion formulation to optimize the generalization capabilities of the generated video.
@@ 3.1 Baseline Transformer Diffusion Network
Baseline Network. The CogVideoX model [yang2024cogvideox] serves as the foundational architecture for our transformer diffusion network, employing a 3D VAE for the compression of video data. In this framework, latent variables are concatenated and reshaped into a sequential format, denoted as $\mathbf{z}_{t}$. Concurrently, the model utilizes the T5 architecture [raffel2023t5] to encode textual inputs into embeddings, represented as $\mathbf{c}_{\text{text}}$. The combined sequences of video latent representations $\mathbf{z}_{t}$ and textual embeddings $\mathbf{c}_{\text{text}}$ are subsequently processed through an expert transformer network. To address discrepancies in feature space between text and video, we implement expert adaptive layer normalization techniques, which facilitate the effective utilization of temporal information and ensure robust alignment between visual and semantic data. Following this integration, a repair mechanism is applied to restore the original latent variable, after which the output is decoded through the 3D causal VAE decoder to reconstruct the video. Furthermore, the incorporation of 3D Rotational Positional Encoding (3D RoPE) [yang2024cogvideox] enhances the model’s capacity to capture inter-frame relationships across the temporal dimension, thereby establishing long-range dependencies within the video framework.
@@
Within diffusion transformers, four primary conditioning mechanisms are identified: in-context conditioning, cross-attention, adaptive layer normalization (adaLN), and adaLN-zero [Peebles2022DiT]. Our investigation primarily focuses on cross-attention and adaptive layer normalization (adaLN). Cross-attention enhances the model’s focus on conditional information by treating condition embeddings as keys and values, while latent representations serve as queries. Although adaLN is effective in simpler conditioning scenarios, it may not be optimal for more complex conditional embeddings that incorporate richer semantic details, such as sequential speech audio. Relevant comparative analyses will be elaborated upon in the experimental section.
@@ 3.2 Audio-Driven Transformer Diffusion
Speech Audio Embedding. To extract salient audio features for our proposed model, we utilize the wav2vec framework developed by Schneider et al. [schneider2019wav2vec]. The audio representation is defined as $\mathbf{c}_{\text{audio}}$. Specifically, we concatenate the audio embeddings generated by the final twelve layers of the wav2vec network, resulting in a comprehensive semantic representation capable of capturing various audio hierarchies. This concatenation emphasizes the significance of phonetic elements, such as pronunciation and prosody, which are crucial as driving signals for character generation. To transform the audio embeddings obtained from the pretrained model into frame-specific representations, we apply three successive linear transformation layers, mathematically expressed as: $\mathbf{c}_{\text{audio}}^{(f)}=\mathcal{L}_{3}\left(\mathcal{L}_{2}\left(\mathcal{L}_{1}\left(\mathbf{c}_{\text{audio}}\right)\right)\right)$, where $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and $\mathcal{L}_{3}$ represent the respective linear transformation functions. This systematic approach ensures that the resulting frame-specific representations effectively encapsulate the nuanced audio features essential for the performance of our model.
@@ 3.3 Identity Consistent Transformer Diffusion
To address this issue, we introduce a control condition within the existing diffusion transformer architecture to ensure long-term consistency of facial identity appearance. We explore four strategies for appearance conditioning: 1) Face attention, where identity features are encoded by the face encoder and combined with a cross-attention module; 2) Face adaptive norm, which integrates features from the face encoder with an adaptive layer normalization technique; 3) Identity reference network, where identity features are captured by a 3D VAE and combined with some transformer layers; and 4) Face attention and Identity reference network, which encodes identity features using both the face encoder and 3D VAE, combining them with self-attention and cross-attention. Our experiments show that the combination with Face attention and Identity reference net achieves the best performance in our model. For further details, please refer to Section 4.3.
@@ 3.4 Training and Inference
(1) Identity Consistency Phase. In this initial phase, we train the model to generate videos with consistent identity. The parameters of the 3D Variational Autoencoder (VAE) and face image encoder remain fixed, while the parameters of the 3D full attention blocks in both the reference and denoising networks, along with the face attention blocks in the denoising network, are updated during training. The model’s input includes a randomly sampled reference image from the training video, a textual prompt, and the face embedding. The textual prompt is generated using MiniCPM[yao2024minicpm], which describes human appearance, actions, and detailed environmental background. The face embedding is extracted via InsightFace[insightface2024]. With these inputs, the model generates a video comprising 49 frames.
@@
Figure 5: Illustration of the dataset, including the flow of data processing, data distribution across different metric, and the visualization of inference data.
@@ 3.5 Dataset
Data Sources The training data used in this work is prepared from three distinct sources to ensure diversity and generalization. Specifically, the sources are: (1) HDTF dataset [zhang2021flow], which contains 8 hours of raw video footage; (2) YouTube data, which consists of 1,200 hours of public raw videos; (3) a large scale movie dataset, which contains film videos of 2,346 hours. Our dataset contains a large scale of human identities and, however, we find that YouTube and movie dataset contains a large amount of noised data. Therefore, we design a data curation pipeline as follows to construct a high-quality and diverse talking dataset, as shown in Figure 5a.
Video Filtering. During the data pre-processing phase, we implement a series of meticulous filtering steps to ensure the quality and applicability of the dataset. The workflow includes three stages: extraction of single-speaker, motion filter and post-processing. Firstly, we select video of single-speaker. This stage aims to clean the video content to solve camera shot, background noise, etc, using existing tools [Plaquet23, Bredin23]. After that, we apply several filtering techniques to ensure the quality of head motion, head pose, camera motion, etc [karaev23cotracker, karaev24cotracker3, Chung16a]. In this stage, we compute all metric scores for each clip, therefore, we can flexibly adjust data screening strategies to satisfy different data requirement of our multiple training stages or strategies. Finally, based on the facial positions detected in previous steps, we crop the videos to a 3:2 aspect ratio to meet the model’s input requirements. We then select a random frame from each video and use InsightFace [ren2023pbidr] to encode the face into embeddings, providing essential facial feature information for the model. Additionally, we extract the audio from the videos and encode it into embeddings using Wav2Vec2 model [baevski2020wav2vec], facilitating the incorporation of audio conditions during model training.
Data Statistics. Following the data cleaning and filtering processes, we conducted a detailed analysis of the final dataset to assess its quality and suitability for the intended modeling tasks. Finally, our training data contains about 134 hours training data, including 6 hours of high-quality data from HDTF dataset, 72 hours YouTube videos, and 56 hours movie videos. Figure 5b also shows other statistics, such as Lip Sync score (Sync-C and Sync-D), face rotation, Face ratio (the ratio of face height to video height).
@@
SadTalker [zhang2022sadtalker] 22.340 203.860 7.885 7.545
DreamTalk [ma2023dreamtalk] 78.147 790.660 6.376 8.364
AniPortrait [wei2024aniportrait] 26.561 234.666 4.015 10.548
Hallo [xu2024hallo] 20.545 173.497 7.750 7.659
@@
SadTalker [zhang2022sadtalker] 50.015 471.163 6.922 7.921 95.194
DreamTalk [ma2023dreamtalk] 109.011 988.539 5.709 8.743 153.450
AniPortrait [wei2024aniportrait] 46.915 477.179 2.853 11.709 88.986
Hallo [xu2024hallo] 44.578 377.117 7.191 7.984 78.495
@@
SadTalker [zhang2022sadtalker] 3.845 10.378 2.953 0.220 470.377 313.758
DreamTalk [ma2023dreamtalk] 4.498 11.005 6.958 1.806 835.480 744.177
AniPortrait [wei2024aniportrait] 1.685 12.025 3.351 1.769 473.173 302.716
Hallo [xu2024hallo] 4.654 10.202 5.268 1.272 394.627 291.052